Operations Report – February 19th

Hi CapCom,

I’m Gabriel, and here is my Operation report.

Crew 189 Operations Report 19Feb2018

SOL: 1

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Roof heater, pipe leak in low deck, ethernet hub, fan of backpack 4

Notes on non-nominal systems:

-Ethernet Hub :

This afternoon an experiment had trouble with an ethernet hub. We’ll try to fix it tonight or tomorrow.

-Fan :

During the engineering check all backpack seemed to be nominal but a fan on backpack 4 turned out of service during the EVA. Tomorrow the crew will be very busy, but we are planning an important backpack fixing operation during the upcoming days.

-Pipe leak :

The flow seems to be reduced, the bucket was empty this morning and at 50% tonight.

-Roof heater :

This morning it was very difficult to control the roof heater. Sometimes the flow became tepid or even cold without any reason because the furnace stops running. I’ve turned on the wall heater which runs all day.

Generator (hours run): All the day (turned on since 6:00 PM 18Feb) because of a cloudy day

Solar— SOC % (Before generator is run at night) 83% at 9 AM

Diesel – 70%

Propane – ?

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons (checked during the EVA).

Trailer to Static Pump used – No but Atila gave a 100 gallons load

Water (loft) – Static to Loft Pump used – YES

Water Meter: 0.24m

Toilet tank emptied: NO

ATV’s Used: None

Deimos rover used: YES

Hours: 109.1

Beginning charge: 100%

Ending charge: 93%

Currently charging: YES

Spirit rover used: NO

Hours: 20.3

Beginning charge: 100%

Ending charge: 100%

Currently charging:

Opportunity rover used: YES

Hours: 15.7

Beginning charge: 100%

Ending charge: 85%

Currently charging: YES

Curiosity rover used: NO

Hours: 12.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: NO

General notes and comments: Very busy today with my own experiments. As soon as possible (from sol 3 I hope) I’ll begin to deal with non-nominal systems (fan, pipe, hub, furnace…).

Questions, concerns and requests to Mission Support: –

Operations Report – February 18th

Crew 189 Operations Report 18Feb2018

SOL: 0

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Pipe leak in low deck, water heater, two thermometers in the Greenhab.

Notes on non-nominal systems: –

Generator (hours run): 14 hours during the night (7 PM to 9 AM)

Solar— SOC % (Before generator is run at night) : 91%

Diesel – ?

Propane – 46.2 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons (estimation based the total height and capacity of the container)

Water (trailer) – 0 gallons.

Water (static) – 400 gallons (estimation based the total height and capacity of the tank)

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter (loft) : 0.44, measurement at 7 PM

Toilet tank emptied: No

ATV’s Used: All ATVs used except the one which doesn’t work (I forgot to check its number)

Oil Added : No

ATV Fuel Used: Gals

# Hours the ATVs were Used today: 20min

Notes on ATVs: –

Deimos rover used: Yes

Hours: 108.7

Beginning charge: ? (not checked)

Ending charge: 98%

Currently charging: No

Spirit rover used: No

Hours: 20.2

Beginning charge: –

Ending charge: –

Currently charging: Yes (Atilia may have unplugged Curiosity to plug in Spirit)

Opportunity rover used: Yes

Hours: 15.3

Beginning charge: ? (not checked)

Ending charge: 93% (100% after being charged)

Currently charging: No

Curiosity rover used: No

Hours: 12.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

HabCar used and why, where : Not used

General notes and comments: Last checks today with Shannon, some things to fix (see Non-nominal systems). Jérémy tried to run his aerosol counter device (LOAC) in the Hab but it doesn’t work. It turns on but doesn’t begin counting. The crew will try to fix it tomorrow.

Summary of internet: –

Summary of suits and radios: All air pipes had been tested and only one doesn’t work (backpack 3 – left one). All radios had been tested and work

Summary of Hab operations: –

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support:

-Thermometers

The GreenHab officer would need a new thermometer.

-Pipe leaks

A bucket is beneath the pipe and receives a liquid which seems to be bad smelling water. Tape would stop the flow, but it only seems to be a temporary solution. Tomorrow I’ll try to find where the leak comes from.

-Water heater

Shannon seems to have fixed it before the very beginning of the SIM. We’ll see it when a crew member will take the first shower of the mission.

-Others:

Some checks have been forgotten (beginning charge of the rovers, numbers of ATVs used…). Sorry for that, I’ll try not to do theses mistakes anymore.

Operations Report – February 17th

SOL: 7

Name of person filing report: Shannon Rupert (for Crew 189)

Non-nominal systems: Power System Control

Notes on non-nominal systems: Nothing new to report

Generator (hours run): Off at 8 am, on at evening (sorry no exact time).

Solar— 79 % SOC

Diesel – not checked %

Propane – not checked psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – not quite 10 gallons. Some were used to fill the tanks for ATV training.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? no

ATV Fuel Used: not recorded Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: All four ARVs were used for training and practice around the campus. Atila took one of the 350’s up Cow Dung Road for samples.

Deimos rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Sojourner rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Opportunity rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes, to town to pick up food and meet crew

General notes and comments: The incoming crew did an awesome job of cleaning the Hab as soon as they arrived.

Summary of internet: Nothing to report
Summary of suits and radios: Nothing to report

Summary of Hab operations: The installation of the stairs, ladders, and railings is complete. While we were going to send photos tonight, the crew would like to wait until tomorrow to send good ones.

I took a shower in the Hab today as my hot water heater is currently broken. I was shocked by the horrible experience. There seemed to be no way to stop the water from getting too hot, then too cold, over and over. I wasted a lot of water just working with the water temperature. After my shower, I turned the hot water heater temperature down to what I think is a good temperature for using the
hot water only (much cooler than the scalding hot temperature that it was on). I will test tomorrow to make sure this is working. If not, we will have to find a solution or we will just waste a crazy amount
of water.

Summary of GreenHab operations: The hydroponics system had a leak and most of the water was lost overnight. We are currently troubleshooting a solution for the project.

Summary of ScienceDome operations: We did more lab work this morning and afternoon working on the halophile research. Eight samples x3 have been plated on 10% NaCl LB.

Summary of RAM operations: Atila will be working in the RAM beginning tomorrow.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing at this time

Operations Report – February 16th

Crew IKEA Operations Report 16 Feb 2018

SOL: 6

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Nothing new to report

Generator (hours run): Off at 8:30 am, on at 6 pm. Turns out the

power system works just fine with lots of welding happening!

Solar— 98% SOC

Diesel – 72 %

Propane – 46 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 10 gallons.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: no

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

Deimos rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 96

Currently charging: Yes

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: Yes

Hours: 20.2

Beginning charge: 100

Ending charge: 75

Currently charging: No

Opportunity rover used: Yes

Hours: 14.9

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes. To town for water for the

aquaponics, packages, searching for the trailer hitch in storage. To

Valley of the Stars for fieldwork. To town to pack food. To town to

get gas for the ATVs and more water.

General notes and comments: How did it get to be Friday already?

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Installations of the stairs is almost

complete! We will have photos tomorrow.

Summary of GreenHab operations: Hydroponics (no fish) set up complete

and water added. We will get it circulating tomorrow. We also

realized that we have two kinds of tomatoes, regular sized and cherry.

Atila and I each had two of the best cherry tomatoes I have ever

eaten.

Summary of ScienceDome operations: We did some lab work this morning.

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Operations Report – February 15th

Crew IKEA Operations Report 15 Feb 2018

SOL: 5

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control
Notes on non-nominal systems: Nothing new to report

Generator (hours run): Generator was once again run all day

Solar— not checked % SOC
Diesel – 72 %
Propane – 47 psi.
Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.
Water (static) – 425 gallons
Trailer to Static Pump used – no
Water (loft) – Static to Loft Pump used – yes
Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: no
Oil Added? n/a
ATV Fuel Used: n/a Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report

Deimos rover used: Yes
Hours:
Beginning charge: 100
Ending charge: 96
Currently charging: Yes

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: Yes
Hours: 20.2
Beginning charge: 100
Ending charge: 75
Currently charging: No

Opportunity rover used: Yes
Hours: 14.9
Beginning charge:
Ending charge:
Currently charging: Yes

Curiosity rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging: No

HabCar used and why, where? Yes, to town for packages. New jumper cables arrived for HabCar. Other set is in Hab.

General notes and comments: Septic tank is almost full. Pumping it out is scheduled for Monday. New battery charger (to replace the one that arrived broken) arrived today and will be stored in the RAM.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: The front airlock door was left unsecured while the contractors were working and the window got blown out again. We will fix it before the crew arrives.

The new stairs are taking shape! Both of the ladders have been installed on the upper deck and the base of the new stairs is in place but not installed. The ladder sections and the lower stairs still need to be installed as well as the railing at the top of the stairs. The contractor expects the installation to be complete on Saturday but the polishing/painting will be done later.

I should be able to send photos tomorrow.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Electrical connection has been scheduled for 3-6 weeks out.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Operations Report – February 14th

Crew IKEA Operations Report 14 Feb 2018

SOL: 4

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Nothing new to report

Generator (hours run): Due to clouds and contractors, the generator was run all day. The guys needed power for welding and sanding, so even when the sun did come out, we decided to just keep it on.

Solar— 99 % SOC

Diesel – 72 %

Propane – 48 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 450 gallons (added 80 gallons)

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: I started and warmed them up today.

Deimos rover used: Yes

Hours: moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: Yes

Hours: Moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes

Hours: Moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes, moved to Hab to pump the remaining water into the static tank. My truck is dead (apparently the battery has gone bad as there is battery acid on top) and so tomorrow we will use the HabCar to go to town and get a new battery for the truck and also pick up the packages that are in town.

General notes and comments: Still muddy most of the day but a strong wind later in the afternoon has dried things out quickly.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: The new stairs and ladders arrived! The contractors began installation and one of the two ladders has been put into place. They are currently bare metal only. We will need to decide which looks best, polished steel or paint. It looks like we will not have time to finish them before the crew arrives on Saturday, so we won’t have to decide right away. Finishing them will probably happen in early May.

Summary of GreenHab operations: Atila continued working on the aquaponics set up and the reorganization of plants. We harvested some greens for salads tonight, and eyed the cucumbers. They may not last until the crew arrives! In addition, we doubled up the pots for the plants in an effort to maintain moisture in the soil longer. In the future, potting soil should be mixed with perlite to ameliorate this problem.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Operations Report – February 13th

Crew IKEA Operations Report 13 Feb 2018

SOL: 3

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: nothing new to report

Generator (hours run): Turned off at 3:30 pm, on at 5:55 pm

Solar— SOC not known %

Diesel – did not check %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 350 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – no

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: none

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

Deimos rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? No

General notes and comments: It rained overnight and was cloudy most of
the day, which was why we kept the generator on most of the day and
did not check some things. Here’s hoping the mud will dry tomorrow.
Everything comes to a standstill in the mud. We did get much work
done on the IKEA project.

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: nothing to report

Summary of ScienceDome operations: nothing to report

Summary of RAMM operations: nothing to report

Summary of health and safety issues: When it is muddy here, it is very
hard to walk. I successfully did go outside and did not fall but I
did not want to go into the buildings and take all the mud in with me.

Questions, concerns and requests to Mission Support: None at this time.

Operations Report – February 12th

Crew IKEA Operations Report 12 Feb 2018

SOL: 2

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: nothing new to report

Generator (hours run): Turned off at 9:30 am, on at 5:00 pm

Solar— SOC 82 %

Diesel – 75 %

Propane – 49 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 350 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: none

Oil Added? n/a

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

Deimos rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 73

Currently charging: yes

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: Yes

Hours: 14.3

Beginning charge: 100

Ending charge: 64

Currently charging: Yes

Curiosity rover used: Yes

Hours: 12.5

Beginning charge: 100

Ending charge: 98

Currently charging: No

HabCar used and why, where? Yes, to town for water

General notes and comments: We continued working on the Ikea project today.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Atila continues his work on the GreenHab and will be sending a GreenHab Report tonight. The grow lights were operating on a different schedule than they should have been on, so we are changing the extra hours on the lights from 7-9 pm to reflect the longer days.

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Okay, mathletes, this is for you:

In the Hab, the kitchen has 111 inches of cabinets and they take up 130 inches of wall space. They are 24 inches deep. The Hab is 30 feet in diameter. Downstairs we want to add 240 inches of cabinets that are also 24 inches deep. How much wall space do we need?

If you are still with me, try this one. The dome is 25 feet in diameter. If we want to put 24 inch deep cabinets on ½ of the wall space, a) how much wall space is that and b) how many inches of cabinetry would we be able to fit in that space?

Extra credit to anyone who can tell us how to calculate countertop dimensions for the kitchen and lab.

We would like our numbers checked. Thank you.

Operations Report – February 11th

Crew IKEA Operations Report 11 Feb 2018

SOL: 1

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Waiting for contractor to troubleshoot

Generator (hours run): Turned off at 9:15 am, on at 6:15 pm

Solar— SOC 92 %

Diesel – did not check %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – no

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: none

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

Deimos rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: not sure

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: not sure

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes, electrical power was restored to the parking
spot on the east side of the GreenHab

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Not sure

HabCar used and why, where? No

General notes and comments: Crew 188 left the Hab at 8:30 am this
morning. They did a nice job of cleaning. We are expecting rain in
the next couple of days so will make another trip for water tomorrow.

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: The Hab is in very good condition. We
spent much of the day working on the Ikea project and have come up
with some exciting changes that will make the kitchen and EVA areas
much easier to use and maintain. We changed the air filter in the
furnace. We are expecting the new stairs, railing and upstairs
ladders to be installed this week.

Summary of GreenHab operations: Atila spent the afternoon working on
the GreenHab. Most of the large plants were dried out and wilting
while others were overwatered. The greens were well tended. We are
limiting the tomato plants to 10, pepper plants to 5 and removing the
melon plants. This is needed in order to have room for the aquaponics
system. There is a cucumber I have my eye on, and if Atila doesn’t
pick it, we will have a nice salad in a couple of days.

Summary of ScienceDome operations: There was new damage to the floor
while Crew 188 was here. There are several small pits and one larger
one in the floor.

Summary of RAM operations: Nothing to report. We are hoping the
electrician can come soon to connect the power.

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: None at this time.

Operations Report – February 9th

Crew 188 Operational Report 09/02/2018

Sol 12

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords, propane, EVA Suits

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 14.5h

Solar: SOC 80%

Diesel: 72%

Propane: 55

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 65 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter: 130949.8

Toilet tank emptied: No

ATVs Used: Yes

Oil Added: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.5

Notes on ATVs: The Gasoline storage area needs to be moved. The sun is creating a large pressure differential on the jerry cans, making that area an explosion risk.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 19.5h

Beginning charge:

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours: 19.1h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 12.0h

Beginning charge:

Ending charge:

Currently charging: No

HabCar: No

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:25

Generator was turned on at 18:24

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: I’ve passed on a report about the EVA suits

Summary of Hab operations: Propane was smelt again this afternoon around 4pm. The heater had been turned off for 24 hours, and nothing else was on. I strongly suggest a licensed contractor investigate the propane leak before another team moves into Hab. There is a significant safety risk as we have not been able to pinpoint the source of the leak.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.